# Universal Narrative Model: an Author-centric Storytelling Framework for Generative AI

### v3

@@ 3.5 Storytelling
The UNM is designed to be human-readable. To make this simple, authors can associate storytelling details with any structural element. For instance, an author may add storytelling details to the Storybeats which roughly correspond to traditional act structure, or dive deeper and describe how each Dynamic, Storypoint, and Storybeat relate to their story. It is entirely up to the author how much storytelling detail they would like to include in each instance of the UNM.
@@ 4 Funding Statement
The Universal Narrative Model was funded by a consortium led by the Entertainment Technology Center at the University of Southern California. The consortium is comprised of Amazon Studios, Disney, Dolby, Epic Studios, and NBCUniversal.
@@ 5 Acknowledgments
The Universal Narrative Model Project team consisted of George Gerba, Paul Bennun, Jim Hull, Brett Leonard, Cecil Castellucci, Chris Huntley, Guy Gadney, and Stephen Stewart.

### v4

@@ 3.5 Storytelling
The NCP is designed to be human-readable. To make this simple, authors can associate storytelling details with any structural element. For instance, an author may add storytelling details to the Storybeats which roughly correspond to traditional act structure, or dive deeper and describe how each Dynamic, Storypoint, and Storybeat relate to their story. It is entirely up to the author how much storytelling detail they would like to include in each instance of the NCP.
@@ 4 Funding Statement
The Narrative Context Protocol was funded by a consortium led by the Entertainment Technology Center at the University of Southern California. The consortium is comprised of Amazon Studios, Disney, Dolby, Epic Studios, and NBCUniversal.
@@ 5 Acknowledgments
The Narrative Context Protocol Project team consisted of George Gerba, Paul Bennun, Jim Hull, Brett Leonard, Cecil Castellucci, Chris Huntley, Guy Gadney, and Stephen Stewart.